OpenStudy (brynndelyn):

Find the non-extraneous solutions of sqrt(x+9)-5=x+4 Choices: x=-2 x=-9 and x=-2 x=9 x=-9

Nnesha (nnesha):

\[\huge\rm \sqrt{x+9} -5 = x+4\] is it ur question ? -5 is not under the rooot tight ?

OpenStudy (anonymous):

I would think so, given the brackets

Nnesha (nnesha):

just wantd to confirm it okay then first add 5 both side and then take square both side

OpenStudy (brynndelyn):

I'm so confused on how to do this problem

Nnesha (nnesha):

\[\huge\rm \sqrt{x+9} -5 = x+4\] is it ur question ? add 5 both side

OpenStudy (brynndelyn):

sqrt{x+9}=x+9

Nnesha (nnesha):

yep now take square both side so you can cancel out square root from left side

OpenStudy (brynndelyn):

would you times each side by 2?

OpenStudy (brynndelyn):

No times by two but raise it to the 2nd power

Nnesha (nnesha):

 like this left side square and square root can cancel out right side (x+9)^2 is same as (x+9)(x+9) so foil (x+9)(x+9)